OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5).

OpenStudy (the_fizicx99):

\[M = \frac{ Y2 -Y1 }{ X2 - X1 }\]

OpenStudy (the_fizicx99):

Plug into the slope formula.

OpenStudy (anonymous):

2/2= 1

OpenStudy (the_fizicx99):

Seriously????

OpenStudy (the_fizicx99):

@e.mccormick

OpenStudy (anonymous):

Answers are: 1 -1 0 1/2

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ -3}})\quad &({\color{red}{ 3}}\quad ,&{\color{blue}{ -5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

I got -2???

OpenStudy (the_fizicx99):

\[M = \frac{ -3 - (-5) }{ 3 - 1 }\]

OpenStudy (jdoe0001):

\(\bf slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ -5}}-{\color{blue}{ -3}}}{{\color{red}{ 3}}-{\color{red}{ 1}}}\implies \cfrac{-5+3}{2}\implies \cfrac{-2}{2}\implies -1\)

OpenStudy (anonymous):

nevermind I did it wrong I got -1? is that right?

OpenStudy (the_fizicx99):

Fffs. Wow, i did the inverse u.u y1 ... nvm

OpenStudy (kewlgeek555):

Just going more in depth to what @tHe_FiZiCx99 said. The slope of a line can be calculated by using the coordinates of any two points on the line. Point one is (x1, y1) and point two is (x2, y2). \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] WARNING:\[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\neq ~Slope\]^Be careful you mix the order up.

OpenStudy (the_fizicx99):

Umm, psst. y1-y2/x1-y2 is possible. y2-y1/x1-x2 isn't.

OpenStudy (the_fizicx99):

Omfg... typo. y1-y2/x1-x2 .... IS possible. Y2 -y1 / x1 - x2 ... Isn't.

OpenStudy (anonymous):

I got the answer guys but thank you!